OpenStudy (anonymous):

a line through (3,1) and parallel to 2x -y =-2 i know you first solve for y so it's y=2x+2

OpenStudy (phi):

so now you know that the slope is 2 (the slope is the number in front of the x in y =2x+2) you can now use point slope form y - y1 = m(x - x1) where m is the slope, and (x1,y1) is a point on the line in your case, m=2 and x1=3 and y1=1 so you get y -1 = 2(x -3) you can use algebra to put this into another form (slope-intercept or standard form, for example) to match the form the answers are in .